A first-order chemical reaction is observed to have a particular reaction rate at 20°C. If its activation energy is 27.5 kJ/mol, how much of an increase in temperature (in °C) would be necessary in order to triple the reaction rate.
Answer #1

Given:
T1 = 20 oC
=(20+273)K
= 293 K
K2/K1 = 3/1
Ea = 27.5 KJ/mol
= 27500 J/mol

use:
ln(K2/K1) = (Ea/R)*(1/T1 - 1/T2)
ln(3/1) = (27500.0/8.314)*(1/293 - 1/T2)
1.0986 = 3307.6738*(1/293 - 1/T2)
T2 = 325 K
= (325-273) oC
= 52 oC

SO,
Increase in temperature = T2 - T1
= 52 oC - 20 oC
= 32 oC
Answer: 32 oC